How large a force is required to accelerate a 1300-kg car from rest to speed of 20m/s in a distance of 80m ?


1
Expert's answer
2022-03-28T11:02:18-0400

The energy-work theorem says

"\frac{mv^2}{2}=Fd"

Hence, the required force

"F=\frac{mv^2}{2d}=\frac{1300*20^2}{2*80}=3250\:\rm N"